OpenStudy (anonymous):

Which value of y makes the equation true? y – 6.3 = –2.2 y =

OpenStudy (anonymous):

watchya think ( ͡° ͜ʖ ͡°)

OpenStudy (kropot72):

y – 6.3 = –2.2 Do you want help to solve this equation?

OpenStudy (anonymous):

not relley but thanks!

OpenStudy (shreenithaa):

y-6.3= -2.2 +6.3 +6.3 y=4.1

OpenStudy (anonymous):

thx